OpenStudy (anonymous):

The table shows prices for three types of clothing. Use a scalar product to find the price with 8.25% sales tax in each item. plain team logo individualize t-shirt $9.00 $13.00 $14.00 shorts $6.00 $9.50 $11.00 jog pants $15.00 $ 21.00 $23.00

OpenStudy (igreen):

Multiply 0.0825 to the values and then add that to the original amount.

OpenStudy (anonymous):

so like $9.00 would be 0.0825+9.00=0.7425+9.00=9.7425

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

No. 0.0825*9.00 + 9.00

OpenStudy (igreen):

No..multiply it

OpenStudy (igreen):

9 + (0.0825 * 9)

OpenStudy (anonymous):

i did i didnt mean to out the addition sighn

OpenStudy (igreen):

Oh..I see..that's correct.

OpenStudy (igreen):

Just do that for every item.

OpenStudy (igreen):

\(\bbox [10pt, lime, border:5pt solid black]{\huge\cal\color{red}\succ\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\prec}}}\) You can give medals by clicking 'Best Response'.